### v1

@@ 6 Conclusions
The diversity of workshops is important to reach multiple pedagogical goals. Large attendance in virtual meetings allows for a global reach and democratization of training and access to computing resources. The size of these virtual meetings was not optimal for spontaneous communication and career network — although future workshops could take this into account. We found that at that scale (200+ participants), it was easier to have participants run containers and software locally, whereas for smaller, virtual meetings, it is possible to provision remote participants with temporary cloud accounts.
In-person meetings are well suited for career development, building collaborations, and provisioning participants with more advanced computing resources, which may be limited to certain countries. These in-person meetings can run longer than virtual events, with the caveat that organizers should pace the delivery of the materials more slowly than they anticipate and even include participant-led hackathons for better learning outcomes and stronger cohort building.
Advanced computing with projects such as JupyterHubs (e.g., Infrastructure-as-a-Service such as Coiled or 2i2c) and science gateways (e.g., Marru et al., 2011; McLennan and Kennell, 2010; Stubbs et al., 2021) promote access to resources, potentially benefiting the user community. Nevertheless, training the community in the concepts of cloud computing and HPC for new developers remains important so that they can continue innovating solutions for large-scale computing for seismological research.
@@
The survey data came from Google Forms responses. Because of the lack of anonymity in the responses, the authors decided not to share the original data. All SCOPED educational materials are open-source (e.g., HPS). Video recordings of our workshops are available on SCOPED YouTube channel.
@@ 7 Declaration of Competing Interests
This work is supported by the Seismic Computational Platform for Empowering Discovery (SCOPED) project under the National Science Foundation (award numbers OAC-2103701 (UW), OAC-2104052(UAF), OAC-2103621 (CSM), OAC-2103741 (CU), OAC-2103494 (UT)). The events were also supported by the eScience Institute, a SCEC grant 22162. The MTUQ workshop was partly sponsored by the Geophysical Detection for Nuclear Proliferation (GDNP) University Affiliated Research Center (UARC) Task Order 7, funded by the Air Force Research Laboratory under contract HQ0034-20-F-0284. The HPS CyberTraining and the SCOPED training were additionally supported by the Southern California Earthquake Center and USGS (SCEC project 22162), by the National Science Foundation (MTMOD, grant no. EAR-2121568, CSA-LCCF, grant no. OAC-2139536, QUAKEWORX, grant no. OAC-2311208), the European Union’s Horizon 2020 research and innovation programme (TEAR ERC Starting; grant no. 852992) and Horizon Europe (ChEESE-2P, grant no. 101093038; DT-GEO, grant no. 101058129; and Geo-INQUIRE, grant no. 101058518).

### v2

@@ 6 Conclusions
The diversity of workshops is essential to reach multiple pedagogical goals. Large attendance in virtual meetings allows for a global reach and democratization of training and access to computing resources. The size of these virtual meetings was not optimal for spontaneous communication and career network — although future workshops could take this into account. We found that at that scale (200+ participants), it was easier to have participants run containers and software locally, whereas, for smaller, virtual meetings, it is possible to provision remote participants with temporary cloud accounts.
In-person meetings are well suited for career development, building collaborations, and provisioning participants with more advanced computing resources, which may be limited to certain countries. These in-person meetings can run longer than virtual events, with the caveat that organizers may consider pacing the delivery of the materials more slowly than they anticipate and even include participant-led hackathons for better learning outcomes and stronger cohort building.
Advanced computing with projects such as services for Jupyter Hubs (e.g., Infrastructure-as-a-Service Iaas such as 2i2c that support centralized servers running Jupyter Lab or Notebooks with multiple-user access), or Python projects that manage distributed cloud resources such as Coiled , and the up-and-coming science gateways (e.g., Marru et al., 2011; McLennan and Kennell, 2010; Stubbs et al., 2021) promote ease of access to resources, potentially benefiting the user community. Nevertheless, training the community in the concepts of cloud computing and HPC for new developers remains important so that they can continue innovating solutions for large-scale computing for seismological research and that their expertise lasts beyond the lifetime of specific Iaas.
@@
The survey data came from Google Forms responses. Because of the lack of anonymity in the responses, the authors decided not to share the original data. All SCOPED educational materials are open-source (e.g., https://seisscoped.org/HPS-book/intro.htmlHPS). Video recordings of our workshops are available on SCOPED YouTube channel.
@@ 7 Declaration of Competing Interests
Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government. We are grateful to associate editor Alan Kafka, reviewers Brad Aagaard and Clara Yoon for their constructive feedback. This work is supported by the Seismic Computational Platform for Empowering Discovery (SCOPED) project under the National Science Foundation (award numbers OAC-2103701 (UW), OAC-2104052(UAF), OAC-2103621 (CSM), OAC-2103741 (CU), OAC-2103494 (UT)). The events were also supported by the eScience Institute, a SCEC grant 22162. The MTUQ workshop was partly sponsored by the Geophysical Detection for Nuclear Proliferation (GDNP) University Affiliated Research Center (UARC) Task Order 7, funded by the Air Force Research Laboratory under contract HQ0034-20-F-0284. The HPS CyberTraining and the SCOPED training were additionally supported by the Southern California Earthquake Center and USGS (SCEC project 22162), by the National Science Foundation (MTMOD, grant no. EAR-2121568, CSA-LCCF, grant no. OAC-2139536, QUAKEWORX, grant no. OAC-2311208), the European Union’s Horizon 2020 research and innovation programme (TEAR ERC Starting; grant no. 852992) and Horizon Europe (ChEESE-2P, grant no. 101093038; DT-GEO, grant no. 101058129; and Geo-INQUIRE, grant no. 101058518). Waveform data, metadata, or data products for this study were accessed through the Northern California Earthquake Data Center (NCEDC), doi:10.7932/NCEDC.
@@
This draft manuscript is distributed solely for purposes of scientific peer review. Its content is deliberative and predecisional, so it must not be disclosed or released by reviewers. Because the manuscript has not yet been approved for publication by the U.S. Geological Survey (USGS), it does not represent any official USGS finding or policy